### v1

# Benchmarking Multi-Agent Preference-based Reinforcement Learning for Human-AI Teaming

Abstract
Preference-based Reinforcement Learning (PbRL) is an active area of research, and has made significant strides in single-agent actor and in observer human-in-the-loop scenarios. However, its application within the co-operative multi-agent RL frameworks, where humans actively participate and express preferences for agent behavior, remains largely uncharted. We consider a two-agent (Human-AI) cooperative setup where both the agents are rewarded according to human’s reward function for the team. However, the agent does not have access to it, and instead, utilizes preference-based queries to elicit its objectives and human’s preferences for the robot in the human-robot team. We introduce the notion of Human-Flexibility, i.e. whether the human partner is amenable to multiple team strategies, with a special case being Specified Orchestration where the human has a single team policy in mind (most constrained case). We propose a suite of domains to study PbRL for Human-AI cooperative setup which explicitly require forced cooperation. Adapting state-of-the-art single-agent PbRL algorithms to our two-agent setting, we conduct a comprehensive benchmarking study across our domain suite. Our findings highlight the challenges associated with high degree of Human-Flexibility and the limited access to the human’s envisioned policy in PbRL for Human-AI cooperation. Notably, we observe that PbRL algorithms exhibit effective performance exclusively in the case of Specified Orchestration which can be seen as an upper bound PbRL performance for future research.
Keywords Multi-Agent  $\cdot$ Human-AI Teaming  $\cdot$ Preference-based Reinforcement Learning
1 Introduction
Modeling co-operation for Human-AI Teaming scenarios has been a long-standing research challenge. With the increase in the pivotal roles AI agents play in day-to-day human activities [1, 2, 3], modeling agents with teaming capabilities from the human-in-the-loop’s perspective has become all the more relevant. This perspective can include what the human-in-the-loop’s expectations are on the AI agent’s policy, what preferences they have on the AI agent’s behavior, or even what their assumptions are on their own policy. Typically, co-operative multi-agent teaming requires all the acting agents to understand each other’s strategies (via explicit communication or implicit deduction from past behavior) and come up with compatible behavior that boosts team performance, and if possible, achieve their personal goals. In this work, we consider the case of a two-agent (Human-AI) cooperative team. While multi-agent cooperation (and even two-agent cooperation) is hard [4], the problem is even more challenging when one of the agents is a human [5, 6, 7]. In particular, it becomes crucial for the AI robot to take into account the cognitive limitations of humans, the preferences humans hold concerning the robot’s goals and strategies, and to align its policies with those of humans to ensure compatibility with policies that humans deem achievable.
Preference-based Reinforcement Learning has been used in several complex domains like robotic manipulation and locomotion to incorporate human preferences on AI agent’s behavior [8]. The initial conceptualization of PbRL focuses on single-agent dynamics with a human in the loop observer available to provide feedback on agent queries [9]. In this work, we provide a first exploration of a more general view of PbRL i.e. in human-AI team setting where the human-in-the-loop co-exists in the same environment as the AI agent and performs actions for achieving the team goal.
The fundamental approach to account for human preferences has been to learn a reward function by querying the human-in-the-loop for their feedback over a pair of agent trajectories. Such preference elicitation becomes more challenging at various levels in a multi-agent scenario. For example, multi-agent (with human-in-the-loop) can span multiple humans or multiple AI agents, which can be cooperative, competitive or act as a swarm system. Moreover, taking human preferences into consideration can be from an external human observing a human-AI team or from the human actively participating in the team, etc. Also, the preferences can themselves be about arbitrary aspects of interaction or behavior between the multiple agents. As a first exploration of Preference-based RL in a multi-agent setting, we constrain our problem to a two-agent cooperative Human-AI team where the human-in-the-loop provides their preferences on the team behavior. Other settings can be explored in future works.
In this work, we specifically look at human-AI teaming as a co-operative game referred to as Human-AI Team PbRL, adopted from [10], with the assumption that the human agent defines the team reward and uses it to compute feasible team strategies, but the AI agent does not have access to it. This reward function is representative of the human’s preferences on the joint behavior of the two agents. We propose a suite of domains, comprising of variants of Highway domain [11] and adapted MuJoCo locomotion domains [12], that requires forced cooperation in a two-agent setting with one agent assumed to be a human agent. As a first exploration of PbRL in Human-AI team setup, we adapt existing state-of-the-art single agent PbRL methods (i.e. RUNE [13], SURF [14] and PEBBLE [15]). Operating under this setting, we argue along two aspects particular of human-AI teaming : Human Flexibility and Access to Human Policy. Human Flexibility, informally, describes how rigid the human agent is with respect to using a specific team strategy. Since the human is acting in the shared world towards a team goal, the robot must find a policy for itself which belongs to set of team strategies considered feasible by the human agent. Otherwise, either the human agent would have to adapt (which can be costlier than having an AI agent adapt) or the team would suffer from sub-optimal performance. Furthermore, the size of this feasible set can be very small or even singleton owing to cognitive limitations of human in computing several strategies or simply due to their preference over team behavior. Another important aspect is the agent’s access to human policy during learning from human preferences. If the agent has zero access to the human’s policy (as selected by the human agent to execute), it has an additional challenge of ‘imagining’ human actions when it queries the human for their preference on the team behavior. On the other hand, the problem can be simplified to a single-agent PbRL setting if the agent has complete knowledge of the human policy and that the human has a single policy in their set of feasible team strategies. We call this as Specified Orchestration, which requires maximal information regarding human behavior and arguably the simplest case for the AI agent, therefore, can be treated as a loose upper bound to the performance of Human-AI PbRL algorithms.
We summarize the contributions of this work as follows:
  1. 1.
    This work serves as a first exploration of Preference-based Reinforcement Learning beyond a single-agent setting.
  2. 2.
    We propose a suite of two-agent domains which require forced cooperation between the two agents to achieve the team objective. The domains are packaged with human’s feasible set of policies (computed through an oracular approach) and can be used for future research in Human-AI team PbRL.
  3. 3.
    We adopt and modify state-of-the-art single-agent PbRL algorithms for our setting and benchmark their performance in policy-learning on the proposed domains.
  4. 4.
    We consider specific aspects of Human-AI team affecting Human-AI PbRL, i.e. Human flexibility in opting for a team strategy and Agent’s access to human Policy, and perform a comprehensive analysis. In particular, we find the case of Specified Orchestration (where the agent has complete access to human’s selected policy and human has a singleton set of team strategies in mind) as a useful upper-bound where the PbRL algorithms exhibit effective performance.
We begin with discussing an example of active human-AI teaming with required co-operation, and analyze the assumptions on the flexibility of the human-in-the-loop in Section 2. Next, we give an overview of the relevant literature in Section 3, followed by describing the preliminaries for this work in Section 4. Section 5 provides a formal framework under which we operate, and defines the notions of Human Flexibility, Agent’s access to Human policy and Specified Orchestration. We describe the proposed domain suite in Section 6.1 and provide an empirical investigation of PbRL for the Human-AI team setup in Section 6. Finally, we conclude with our discussion on the learnings from this work motivating PbRL for human-AI teaming problems in Section 7. An appendix has also been attached as supplementary that includes additional details on the experiments and results, for the ease of reproducibility of our work for future research.
2 Active Human-AI Teaming Example
Figure 1: Two Arm Peg-In-Hole Environment: A representative example for Human-AI Team PbRL Cooperation Game.
Let us consider the example of the Baxter robot, a classic bi-manual robot as shown in Figure 1. There are two robotic arms where one is holding a peg while the other is holding a square plate with a hole in it. Assume that the arm holding the peg is being controlled by the human agent. The human agent has already decided on a policy that they will be moving the peg to a specified location, and insert the peg according to their policy. While there may exist multiple corresponding robot policies that successfully allow for team success, the human agent has specific preferences on the robot behavior as well. For example, the human wants the robot to hold the plate horizontally, reach the location where the peg would be in time and the robot arm should slowly approach that location for safety. The joint task for the two agents is to place the peg in the hole (square plate) successfully without collisions in accordance with the human preference.
The agent is unaware of its goals (holding the plate horizontal at some location), the human preference on how to achieve that goal (manner of reaching the location) and the human partner policy (where the location of peg would be and the trajectory of the human controlled arm to avoid collisions). Therefore, it must utilize some form of communication to gather information about these aspects. Such communication may happen through explicit language-based interaction (which may require an expert understanding of the internal working of the robot, and a robust method for reward function specification that is not prone to reward-hacking), through demonstrations (but in this scenario the human agent holds a peg whereas the robot agent holds the plate which makes providing a demonstration to teach the robot non-trivial) or through binary feedback on agent’s queries (PbRL). In this work we explore the latter case of studying PbRL, where the agent generates behavior trajectories with imagined actions of the human and itself and queries a pair of these behaviors for human feedback. The human agent views the queried trajectory (typically presented as a video) pair and specifies which they prefer more, i.e. which pair achieves the team objective according to them. Recall that the human agent has pre-computed a policy for themselves and has an expectation from the AI partner to complete any remainder objectives, therefore the preference labels are given according to how well the AI-agent is conforming to the human-expected policy. Naturally, the key challenges for the AI agent are to realize the human partner’s chosen policy (such that the queried behaviors show reasonable imagined human agent actions), and how large the team-behavior solution set is based on the human’s internal reward function (for example, the human may find it acceptable to have the plate be placed somewhere in a large region at any angle).
3 Related Work
Preference-based Reinforcement Learning: Preference-based RL has become a popular paradigm for human-in-the-loop settings where the human is queried for binary feedback by the AI agent to learn a reward function [16, 17, 18, 19, 20, 21, 22]. These methods have not only shown to be effective in teaching an AI agent to accomplish a task where rewards may be otherwise hard to specify such as in the case of tacit knowledge tasks, but also to incorporate the human-in-the-loop’s preferences on how the task needs to be achieved [17]. However, PbRL methods have only been seen in single-agent settings where the AI agent is acting in the environment, and the human is only interacting with the system as and when asked for query feedback [23, 24, 25, 26]. Having realized the potential of PbRL frameworks, we see its applicability in human-AI teaming scenarios where both the human and the AI agent act in the environment to accomplish a team task, and the human may want to specify their preferences through binary feedback distilled into the agent’s reward function. Closest to our research are prior works in multi-agent Inverse Reinforcement Learning (IRL) [27, 28, 29]. We note that IRL and PbRL have key distinctions in assuming expertise of the human in the loop and the modality of the feedback. [30] were one of the first to tackle the multi-agent IRL problem with a central controller. Several works have attempted to study multi-agent IRL for homogeneous agent systems [31, 32, 33, 34] whereas our focus is on a heterogeneous system with a two-player Human-AI cooperative team. More recently, [10] proposed the Cooperative Inverse RL formulation to highlight the limitation of classic formulation which does not allow humans to provide teaching demonstrations. In contrast, our focus is on allowing the AI agent to make explicit queries to the human agent to elicit preferred behaviors. Further, while IRL based setups are dependent on expert-demonstrations by the human agent, such demonstrations may be infeasible (in our Peg-In-Hole example, the human does not control the arm with the square plate), or the human maybe incapable. PbRL, on the other hand, makes it easier for the human agent to provide a binary feedback over trajectories generated by the agent.
Multi-Agent Teaming : Deep Reinforcement Learning has also been used in multi-agent settings [35, 36, 37] such as non-competitive games [38, 39], zero-sum [34], general-sum games [33, 40] and other collaborative environments where rewards are not shared across agents [41, 42]. More recently several works have investigated the challenges in Human-AI teams but commonly assume robot’s access to the team reward function [6, 43, 44, 45]. Prior works have also studied joint-decision making in a two-agent collaboration setting [46, 47, 48, 49] but do not consider inverse RL or PbRL setting.
4 Preliminaries
Preference-based Reinforcement Learning : In traditional PbRL setups, an agent $A$ interacts with the environment $\mathcal{E}$ with a missing reward function $\mathcal{R}$. It assumes the underlying problem as a MDP $\mathcal{M}=\langle\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R},\gamma\rangle$. The agent $A$ may not have access to this reward system $\mathcal{R}$ as it may either be hidden [50], hard to specify as seen in tacit-control tasks [17], or otherwise may not be able to capture the goal completely as seen in human-in-the-loop scenarios [51].
The problem of learning the unknown reward function $r_{\psi}$, in single-agent settings, is formulated as a supervised learning problem [17]. A classifier is trained on limited human feedback over agent trajectories. Formally, each agent trajectory is a segment $\sigma$, a sequence of agent observations and actions $\{(s_{k},a_{k}),(s_{k+1},a_{k+1}),\cdot\cdot\cdot(s_{k+N-1},a_{k+N-1})\}$ where $N$ represents the horizon. The human-in-the-loop, i.e., the teacher, gives a feedback $y\in\{0,1\}$ to the agent when queried with a pair of trajectories ($\sigma_{1},\sigma_{2}$), and this feedback is stored as a label along with the two segments in a trajectory buffer $\mathcal{D}$ as ($\sigma_{1},\sigma_{2},y$). Following the Bradley-Terry preference model [52], the preference predictor using the reward function $r_{\psi}$ is learned:
$P_{\psi}[\sigma_{0}\succ\sigma_{1}]=\frac{\exp\Bigl{(}\sum_{t}\hat{r_{\psi}}(s_{t}^{0},a_{t}^{0})\Bigl{)}}{\sum_{i\in\{0,1\}}\exp\Bigl{(}\sum_{t}\hat{r_{\psi}}(s_{t}^{i},a_{t}^{i})\Bigl{)}}$ (1)
where, $[\sigma_{0}\succ\sigma_{1}]$ implies that trajectory segment $\sigma_{0}$ is preferred over trajectory segment $\sigma_{1}$ by the human teacher. Given the trajectory buffer $\mathcal{D}$, the reward function $r_{\psi}$ is learned by training on minimizing the following Cross-Entropy loss:
$\mathcal{L}_{CE}=-\displaystyle\mathop{\mathbb{E}}_{\begin{subarray}{c}(\sigma_{0},\sigma_{1},y)\sim\mathcal{D}\end{subarray}}\Bigr{[}y(0)\text{log}P_{\psi}[\sigma_{0}\succ\sigma_{1}]+y(1)\text{log}P_{\psi}[\sigma_{1}\succ\sigma_{0}]\Bigr{]}$ (2)
5 Preference-based RL for Human-AI Teaming
Preference based RL has been predominantly studied in single-actor human-observer setting where the human provides feedback on the agent’s behavior but does not act in the shared environment. The aim of this work is to explore PbRL for human-AI team scenario. As discussed in Section 1, there can be multiple interpretations of multi-agent (or a two-agent human-AI team) PbRL. We consider the setup where the human and the AI agents are a cooperative team and the role of the robot is to achieve goals as set by the human agent. Having no knowledge of its objectives, the robot relies on a reward learning process through Preference-based Reinforcement Learning to elicit and learn about the human preferences over the team behavior. We formalize our Human-AI Team PbRL as a special case of the Cooperative Inverse Reinforcement Learning formulation [10] in section 5.
Further, we discuss the two critical aspects for Human-AI team i.e. Human Flexibility over team behavior 5.2 and assumption on agent’s access to human policy 5.3. We discuss a special case with assumed knowledge about human’s policy and the fact that human admits a single team behavior as an optimal solution, referred to as Specified Orchestration. Since we adapt PbRL methods developed for single-actor human-observer setups for our first exploration of Human-AI team PbRL, we discuss the design decisions in Section 6.2.
5.1 Human-AI PbRL Cooperation Game
[10] introduces the problem of Cooperative Inverse Reinforcement Learning but studies it for Apprenticeship Learning. For our Human-AI Team PbRL setting, we modify their formulation as follows:
Definition 5.1.
A Human-AI PbRL cooperation game, as a special case of a Cooperative Inverse Reinforcement Learning game, $M$ is a two-player Markov game with identical payoffs between a human $H$ and a robot $R$. The game is defined as a tuple $M=(S,\{{\mathcal{A}^{H},\mathcal{A}^{R}}\},T(\cdot|\cdot,\cdot,\cdot),R_{\theta}(\cdot,\cdot,\cdot),\gamma)$ with the following definitions:
  • $S$ a set of world states: $s\in S$;
  • $\mathcal{A}^{H}$ a set of actions for H : $a^{H}\in\mathcal{A}^{H}$;
  • $\mathcal{A}^{R}$ a set of actions for R : $a^{R}\in\mathcal{A}^{R}$;
  • $T(\cdot|\cdot,\cdot,\cdot)$ a conditional distribution on the next world state, given previous state and action for both agents, i.e.,
    $T(s^{\prime}|s,a^{H},a^{R})$;
  • $R_{\theta}(\cdot,\cdot,\cdot)$ a reward function that maps world states, joint actions to reward values according to H’s preference on the team objective, H’s objectives, R’s objectives and preference over R’s behavior;
  • $\gamma$ a discount factor : $\gamma\in[0,1]$
The reward function $R$ is set by the human agent defining the team objectives and human-agent acceptable policy pairs $(\pi_{H},\pi_{R})$. The game begins at the initial state $s_{o}\in S$. The human agent pre-computes acceptable solutions to the game with assumed knowledge on AI partner’s capabilities. Suppose the computed solution set is $\Pi=(\pi_{H},\pi_{R})_{0},\pi_{H},\pi_{R})_{1}\cdots\pi_{H},\pi_{R})_{n}$. For simplicity, suppose the solution set is singleton. The human takes an action following their policy $\pi_{H}$. The robot takes a random action in $\mathcal{A}^{R}$. As the game proceeds the robot builds a bank of trajectories from past experiences. Following PbRL setup, the agent samples two trajectories from this bank and queries the human agent for their feedback. The human agent uses their internal reward function $R$ to provide their preference label. A behavior pair where the shown human actions align with the true human actions from $\pi_{H}$ and the shown robot actions align with human expected $\pi_{R}$ such that resultant return of the trajectory is higher than the alternate is preferred.
While the policies can be arbitrary functions of history to next action as $\pi_{H}:[\mathcal{A}^{H}\times\mathcal{A}^{R}\times S]*\rightarrow\mathcal{A}^{H}$ (and similarly for robot R), we restrict the policy to be defined on states (thereby assuming the dynamics, reward function and the optimal policy pair is Markovian with respect to the world state).
5.2 Flexibility of the Human-in-the-Loop
It is important to note that for our human-AI teaming PbRL setup, we only consider the case where the AI agent’s access to the underlying reward function is restricted. That is, the agent does not experience a sampled reward from the environment, and instead, depends on an explicit reward learning procedure (through human feedback) to realize its objectives in the team. Assume the set $\Pi=\{\Pi_{1},\Pi_{2},\cdots\Pi_{n}\}$ where $\Pi_{i}=(\pi_{i}^{H},\pi_{i}^{R})$ be the possible solutions of the Cooperation Game defined in Section 5.1 as can be computed by the human. We define human flexibility over their policy, i.e. $\{\pi_{i}^{H}\}_{\{i=1\cdots n\}}$ as :
Definition 5.2.
Given $\Pi$ solution set, the human in the loop is said to be $\epsilon$-H-flexible if
|$\left|\{\pi_{k}^{H}\,|\,k=1\,\text{to}\,n\}\right|=\epsilon-1$
The definition of H-flexible (5.2) human in the loop helps us recognize the rigidity of the human agent over a strategy for the two agents. A large $\epsilon\rightarrow|\Pi|$ implies that the human is willing to execute any one of the policies in $\Pi_{H}$ subsequently providing the robot to converge to a larger solution space (policies $\pi_{k}^{R}$). The nature and requirements of the cooperation for the given team-goal can be helpful in determining the values of $\epsilon$. For instance, $\epsilon=0$ is useful when considering human cognitive limitations, thereby the human is able to come up with a single policy for themselves.
When $\epsilon>1$ then the human agent has a choice to make in selecting which policy $\pi_{k}^{H}$ they will follow. In this work, we assume that the AI agent’s objective is to be conformant to all the human policies. This choice is motivated from how prior works in Human-AI cooperation have attempted to learn the robot’s policy [53]. Other alternatives are: the agent is compatible to at least one of the human policies, a subset of the human policies etc. exist which would also impact the training regime of the RL agent (i.e. which policy is used by the human agent when the Robot is learning its policy?). Following prior works [53, 54] in learning a response policy, during training we randomly sample the human’s policy which is used for an entire episode.
5.3 Access to Human Policy
Based on the Human-AI Team PbRL cooperative game describe in Section 5.1, the human and the robot take actions in a shared environment. This implies that for the duration of training, in addition to the preference feedback, the framework assumes the human to sample their optimal action $a_{h}$ for a world state $s$. That is, the agent gains indirect access to the human policy by virtue of the shared world state and the fact that human-AI coexist in the shared world. While this is the default operating assumption through this work, it has higher requirements from the human agent as compared to single-agent PbRL. For example, in single-agent PbRL the human is required to be present in the loop during agent’s training procedure and provide feedback, whereas in the current setup, in addition to the preference feedback the human is also required to follow their policy throughout the training process (which can be millions of steps).
We can reinterpret the Human-AI Team PbRL cooperative game as follows: the agent has access to an environment simulator which can simulate the agent’s actions and the human’s actions. Now, the agent uses the policy $\pi_{R}$ being learned and an imagined human policy $\pi_{R}^{h}$ to generate team behaviors. These behaviors can be stored in a trajectory buffer and queried to the human actor. The advantage of this re-interpretation is that while the human agent is present in the loop they are not required to execute millions of actions during agent’s training. The disadvantage is, the agent has to model the human’s policy in addition to the human’s preferences. This re-interpretation of the cooperative game can enable us to study the performance of the PbRL agent when it has budgeted access to human actions. Full access is equivalent to the default interpretation, 50% access is when the agent can query for the human’s actions for the first half of the training process and 0% access implies the agent cannot query human’s actions at all. In this work, the agent defaults to using random actions as imagined human actions when it exhausts the human-action query budget. Techniques like behavior-cloning [6] can be leveraged to independently model the human’s policy along with the PbRL step for modeling human preferences in future work.
5.4 Specified Orchestration
Recall our Two Arm Peg-In-Hole example in Section 2. It is clear that since the human agent defines the team objective through reward function $R$, they are restricting the possible robot behaviors to achieve the team goal. For example, the robot arm should approach very slowly towards the human controlled arm because of human’s preference for safety. It becomes extremely challenging if the agent does not have access to the human’s policy. Further, as described in Section 5.2, if the human is switching their policy in the feasible team behavior set during the agent’s training process, the agent now has an additional objective of learning a conformant response policy that is compatible with all of the policies in $\epsilon$-Human Flexible solution set. Our empirical results in sections 6 show that the agent’s performance in policy learning is significantly diminished in the absence of these assumptions.
Definition 5.3.
Specified Orchestration is a special case of the Human-AI Team Cooperative game $M=(S,\mathcal{A},\mathcal{T},\mathcal{R},\gamma)$ where the human agent (H) is 1-H-flexible with the team behavior solution $(\pi_{H},\pi_{R})$ and $\pi_{H}$ is accessible by robot R.
Note that this is a special case of the cooperative game where we have specific assumptions on the human’s preference on the team behavior and access to human’s actions. While these assumptions are quite expensive to meet, they play an important role in benchmarking Human-AI Team PbRL algorithms. Specified Orchestration provides a loose upper-bound performance by assuming away other challenging aspects of Human-AI interaction except preference elicitation, reward learning and subsequent policy learning.
6 Benchmarking Human-AI Team PbRL
In this section we will discuss the proposed benchmark suite (Section 6.1) and our experiments showcasing performance of Human-AI Team PbRL across the two critical aspects (as discussed in Sections 6.4 and 6.3) i.e. impact of human flexibility on team behavior and the impact of providing access to human policy to the PbRL agent.
Further, our discussion will highlight the challenges associated with using PbRL for Human-AI team cooperative game and empirically answer the following questions :
  1. 1.
    What is the relative performance when the human partner can choose from multiple possible policies to follow, for accomplishing the team task?
  2. 2.
    In Human-AI team PbRL, what is the performance of the adapted single-agent PbRL algorithms with maximal assumptions on access to human policy and $\epsilon=1$-H-flexible human (Specified Orchestration)?
  3. 3.
    How do PbRL algorithms perform with a budgeted access and no access to human’s policy?
Question (1) supports our discussion on Human Flexibility, (2) highlights the case where PbRL algorithms are effective making Specified Orchestration a loose upper bound performance for studying Human-AI team PbRL performance. Question (3) discusses the special case of limiting agent’s access to human policy thereby highlighting the need for an explicit modeling of human policy.
In the subsequent sections we first study (Q2) which provides us with an upper bound performance of PbRL algorithms followed by (Q3) where we find that the adapted PbRL algorithms fail without access to human policy. Finally, we study (Q1) under full-access to human policy with a $\epsilon>1$-H-flexible solution set.
We refer the readers to the supplementary material (Sections A, B, and C) for detailed discussion on algorithms, the human policy and preference oracles, domain construction and the respective hyperparameters used for the experiments.
Figure 2: MA Highway domain - human and the AI agent cars are shown one behind the other.
Figure 3: MA MuJoCo domains - (from L to R) Cheetah, Ant, Walker, Swimmer, & Hopper. The action space for the joints in each of the domains has been split between the human and the AI agent.
Algorithm 1 Adapting Single Agent PbRL for Human-AI Team
  Input: feedback frequency $K$, # queries per feedback session $M$, human policy set: $|\Pi|$, agent access budget: $b$
  Initialize parameters of $Q_{\theta}$ and $\hat{r}_{\psi}$
  Initialize a dataset of preferences $D_{h}\leftarrow\varnothing$
  // EXPLORATION PHASE
  $D_{\tau},\pi_{\phi}\leftarrow$ EXPLORE() in [15]
  POLICY LEARNING PHASE
  for each iteration do
     // REWARD LEARNING PHASE
     if new episode then
        choose human policy $\pi_{H}\in|\Pi|$
     end if
     if iteration % $K$ == 0 then
        for m in 1…$M$ do
           $(\sigma^{0},\sigma^{1})\sim$ SAMPLE() in [15]
           Query human partner for $y$
        end for
        for each gradient step do
           Sample mini-batch $\{{(\sigma^{0},\sigma^{1},y)}_{j}\}^{D_{h}}_{j=1}\sim D_{h}$
        end for
        Relabel entire replay buffer $D_{\tau}$ using $\hat{r}_{\psi}$
     end if
     REINFORCEMENT LEARNING PHASE
     for each time-step $t$ do
        if time-step $t\leq$ budget $b$ then
           sample $a_{h}\sim\pi_{H}$
        else
           randomly sample $a_{h}\sim\mathcal{A}^{H}$
        end if
        Collect $s_{t+1}$ by taking $a_{t}\sim\pi_{\phi}(a_{t}|s_{t})$ and $a_{H}$
        Store transitions $D_{\tau}\leftarrow D_{\tau}\{(s_{t},a_{t},{\color[rgb]{1,0,0}a_{h}},s_{t+1},\hat{r}_{\psi}(s_{t}))\}$
     end for
     for each gradient step do
        Sample random mini-batch $\{{\tau_{j}}\}^{D_{\tau}}_{j=1}\sim D_{\tau}$
        Optimize $L^{SAC}_{critic}$ and $L^{SAC}_{actor}$ w.r.t. $\theta$ and $\phi$, respectively [15]
     end for
  end for
6.1 Evaluation Domains
6.1.1 Multi-Agent Highway Orchestration Task:
We first design a set of three multi-agent environments built on top of the Highway environment [11]. As shown in Figure 2, we construct a three lane highway where the human car (in orange) is on the top-most lane, and the AI agent (in purple) has to overtake and merge back into the human car’s lane by increasing its speed and avoiding collisions with any other car in the scene. The human prefers that the AI agent car overtakes from the right-most (bottom-most in the figure) lane and not from the middle lane. Similarly, we construct two more scenarios. In the second one, both the cars are in the right lane and the human preference is for the AI agent car to overtake from the left-most lane. In the third domain, both cars are in the middle lane and the AI agent can only overtake from the left lane. In the results shown in Figure 4, these environments are referred to as MA Highway - Right Overtake, MA Highway - Left Overtake, and MA Highway - Middle Overtake respectively.
6.1.2 Multi-Agent Locomotion Orchestration Tasks:
We present a more complex domain suite, which necessitates cooperation, to study Human-AI team PbRL cooperation game. Following prior works in constructing multi-agent environments [55] in continuous control domains, we modify the MuJoCo-based locomotion environments [56] for the human-AI teaming tasks. The locomotion tasks include a robot attempting to move (for example by a walk, run or crawl). Such tasks are essentially tacit knowledge robotic simulation tasks for which reward specification can be challenging [14, 9]. Moreover, single-agent variant of these environments have been used as benchmark domains by single-agent PbRL works [14, 13, 15], further motivating their use.
We construct multi-agent settings for Cheetah, Ant, Walker, Swimmer & Hopper environments, as shown in Figure 3. In each environment, the human agent controls the orange highlighted joints action space, while the rest of the joint angle controls are for the AI agent highlighted in purple. For example, in the MA-Walker, the human controls the walker’s right leg joints to walk forward, while the AI agent controls the left leg joints.
Figure 4: Learning curves on MA Highway (rows 1 to 3) and MA MuJoCo - Cheetah (row 4): comparing (from L to R) (a) 0% access to $\pi_{H}$, (b) partial access to $\pi_{H}$, and (c) human flexibility on multiple $\pi_{H}$, with Specified Orchestration case that assumes a single $\pi_{H}$ and complete access to it; as measured on the episodic returns. The solid lines and shaded regions represent the mean and standard deviation, respectively, across three runs.
((a))
((b))
((c))
((d))
Figure 5: Learning curves on MA MuJoCo domains: comparing PbRL algorithms extended to the multi-agent setting under Specified Orchestration, as measured on the episodic returns. The solid lines and shaded regions represent the mean and standard deviation, respectively, across three runs.
6.2 Adapting Single Agent PbRL to Human-AI Team setting
We present our modified PbRL algorithm for the human-AI team setting, which uses the PEBBLE algorithm [15] as a backbone. Note, we make the same modifications for experiments using RUNE [13] and SURF [14] as these algorithms also use PEBBLE as their PbRL backbone. We do not require any change to the model architectures (including reward approximator and policy approximator). The updates are highlighted in red (see Algorithm 1) and can be easily incorporated into any online single agent PbRL algorithm.
Specifically, we assume a human policy set $|\Pi|$ from which the human agent can select a single (or multiple) policies for themselves, and the AI agent has the access to the human policy during training with a budget $b$. When a new episode begins, we uniformly sample a human policy $\pi_{H}$ from $|\Pi|$. Note, if the human only has a single policy, i.e., in the case of Specified Orchestration, then we use the same $\pi_{H}$ for every episode. During the policy learning phase for the AI agent if the agent has remaining budget to request actions from human policy, we sample $a_{H}$ from $\pi_{H}$, or else, the AI agent assumes a random action on behalf of the human. Again, in the case of Specified Orchestration, the budget $b$ allows the AI agent to access human actions till the end of the training. As per our re-interpretation of the Human-AI team PbRL Cooperation game in Section 5.1, this is equivalent to the assumption that human is acting in the shared world with the agent.
Following past works in Preference based Reinforcement Learning, we use a simulated oracular human policy to provide the actions on behalf of the human agent. As suggested in [22] we use a scripted preference labeling function that provides preference labels on behalf of the human agent to the PbRL agent. We train SAC [57] by assuming centralized controller for human and AI agent using underlying task reward available with the environment package for MuJoCo [12] and handcrafted reward function for Highway. We take the actions corresponding to the joint angles under control of the human as human actions and rest for the AI agent’s actions expected by the human as part of the compatible policy pair. Preference labels are generated based on hand-crafted rewards for MA Highway and environment packaged rewards (which serve as team reward) for MA MuJoCo. We refer the readers to the appendix for further details.
6.3 Impact of Access to Human Policy
6.3.1 Specified Orchestration - complete access to $\pi_{H}$:
To answer Question (2) and to support our discussion on Specified Orchestration, we report the performance of all three PbRL algorithms on the three MA Highway, and five MA MuJoCo environments in Figures 4 and 5. We note from each of the plots, that PEBBLE [15], RUNE [13] and SURF [14] show the optimal returns. From Figure 5, we also note that RUNE consistently performs better than SURF and PEBBLE in these multi-agent settings.
6.3.2 Restricting Access to Human Policy:
To answer Question (3), we first investigate the case where we restrict the AI agent’s budget to access the human policy $\pi_{H}$ to 50% for MA Highway domains, and 10% for MA MuJoCo - Cheetah. We allocate this entire budget in the first half of the training. For example, in the case of MA MuJoCo - Cheetah domain, the AI agent has access to $\pi_{H}$ up-till 100k steps, and assumes random actions sampled for the remaining training steps. We report the results in the second column of Figure 4. Surprisingly, we note that the performance plummets significantly as soon as the human agent’s actions are not available for all the domains.
6.3.3 No Access to Human policy:
In reference to Question (3), we report the performance of PEBBLE [15], RUNE [13], and SURF [14] on the three MA Highway and MA MuJoCo - Cheetah domains in first column of Figure 4. None of the algorithms have any access to $\pi_{H}$, and thus, have to ‘imagine’, i.e., assume random actions for the human agent. In all the four domains and three algorithms, we note that the performance is significantly worse as compared to the case when the AI agent has complete access to $\pi_{H}$. Utilizing explicit modeling of human policy is a natural next step and an important future work, but it beyond the scope of the first exploration of PbRL in a Multi-agent setting.
6.3.4 Observed Challenges:
From the experiments with 0% and partially restricted access to the human partner policy, we observe that reward learning is severely impacted. Since Specified Orchestration assumes maximal information on the human partner’s policy and preferences, the team performance is optimal. This highlights the potential of using Specified Orchestration as a means to obtain an loose upper bound on the Human-AI team PbRL cooperation game performance. Further, unless the AI agent explicitly models human policy (and preferences) with a restricted access, reward learning along with accomplishing the team task stays a challenging problem.
6.4 Impact of Human Flexibility
6.4.1 Multiple Human policies:
We observe that each of three PbRL approaches perform the best in the case where the agent had complete access to $\pi_{H}$. To study the impact of Human Flexibility (our Q1) on PbRL performance we operate under the assumption that agent has complete access to the human policy (i.e. human executes their policy during agent’s training)
We assume that for each episode of learning, the human agent can sample 1-out-of-N (3, in our case) policies to provide the actions to the AI agent. From the results shown in the third column of Figure 4, we note that the performance comparatively drops as in the case where we assumed a single $\pi_{H}$ and complete access to it. The decline in performance is amplified in the more complex case of MA MuJoCo Cheetah where cooperation is forced amongst the two agents. We conclude that the team performance is the highest in the case of Specified Orchestration, where the human agent has a single $\pi_{H}$ and the AI agent has complete access to the same.
6.4.2 Observed Challenges:
From the experiments with multiple $\pi_{H}$ available to the AI agent during the training phase, we conclude what while it is a cognitively challenging task for a human agent to compute multiple policies and also provide feedback in this case, reward learning is also negatively impacted as seen from our empirical analysis.
7 Conclusion & Future Work
We perform a first exploration of Preference-based Reinforcement Learning for Human-AI team setup. We introduce a Human-AI PbRL Cooperation game as a special case of Cooperative Inverse Reinforcement Learning [10] and argue across two critical dimensions, namely - AI agent’s Access to Human Policy and Human Flexibility. In our heterogeneous Human-AI cooperative setup, both the agents are rewarded according to human’s reward function for the team comprising of human’s objectives for the team, sub-objectives for themselves and the robot, and human preferences on team behavior. For investigating the performance of PbRL approaches, we modify the existing state-of-the-art single-agent PbRL algorithms (namely, PEBBLE [15], SURF [14] & RUNE [13]). To this end, we benchmark PbRL algorithms on our benchmark suite consisting of simpler Highway domains and more complex Robot locomotion.
We note that all three PbRL algorithms gave optimal returns only when the human partner has a preference for a single policy pair (team behavior) and that they have complete access to human partner’s policy (termed as Specified Orchestration). We discuss how Specified Orchestration can be useful to benchmark future works in Human-AI team PbRL as a loose upper bound on algorithm performance. To our surprise, partial access to $\pi_{H}$ also had a drop in performance as soon as the agent’s access budget was exhausted. Subsequently, we also observe poor performance when the AI agent does not have access to human partner’s policy. We conclude that there is a need for AI agents to explicitly model partner human policy (along with their preferences). Furthermore, utilizing these observations, we assumed complete access to $\pi_{H}$, and ablated in the case where the human partner is flexible, i.e., has multiple acceptable team strategies. Counter-intuitively, we find that when the human is more flexible, the agent’s performance declines. This is primarily because we constrain our AI agent to learn a conformant policy (compatible with all human policies).
From our benchmarking results, we conclude that Specified Orchestration, while assuming the maximal amount of information on the human, shows notable performance across all domains. However, this can also be seen as one of the key limitations as it may not always be the case that the human can easily compute and provide their policy to the AI agent (or be available to act during the training procedure of the AI agent). Secondly, while this approach allows for optimal team performance of an AI agent with a specific human, it may be hard to generalize the learnt reward function (and hence, the learnt policy) to multiple human partners. Lastly, as in the case of single-agent PbRL works where human is just an observer, the problem of expensive human interaction with the AI agent is carried over to our setting. To conclude, our work provides researchers with a first exploration of PbRL beyond single-agent setup, towards human-AI teaming.
8 Societal Impact
This paper focuses on motivating the use of Preference-based Reinforcement Learning (PbRL) for human-AI teaming settings. We argue on two critical aspects of such teaming settings in this work, namely - Human Flexibility and AI agent’s Access to Human Policy. The empirical benchmarking includes experiments across a suite of domains with simulated human oracles. Since this work uses simulated human oracle policies and preferences for the experiments, it does not bear any direct ethical implications. However, we note that once such studies advance to real-world teaming problems, ethical considerations with respect to privacy, adherence to the intended preference, consent, transparency, etc. will come in place.
Specifically, with the use of PbRL methods in real-world settings, we highlight that the human partner may experience unexpected and unwanted scenarios while interacting with the AI agent. This may include dealing with irrational or erratic behavior of the AI agent. This may further lead to frustration and fatigue for the human which may be undesirable. Moreover, trust on the AI agent can also play a crucial role in the effectiveness of modeling a human-AI teaming problem. For example, if the human partner does not trust the AI agent for task completion or its ability to learn their preferences (due to any possible reason), the human may find it difficult to give feedback that can be useful for the AI agent’s training. Hence, generalizing PbRL for real-world human-AI teaming settings may require design decisions that can encompass these aspects to benefit the human partner agent.
Appendix A PbRL Algorithm
A.1 PEBBLE Algorithm
PEBBLE is a PbRL algorithm that comprises of two key elements: pre-training and relabeling experience buffer. To gather a wide range of experiences, PEBBLE starts by using intrinsic motivation [58, 59, 60, 61] to pre-train the policy, which optimizes the policy to increase the state entropy in order to explore the environment effectively. Afterwards, PEBBLE uses the SAC algorithm, a state-of-the-art off-policy RL algorithm, to further train the policy. To ensure stability in the learning process, PEBBLE relabels all experiences in the buffer when the reward model is updated.
A.2 Modified PEBBLE Algorithm for varying Access to Human Policy experiments
Algorithm 2 Adapting Single Agent PbRL for Human-AI Team
  Input: feedback frequency $K$, # queries per feedback session $M$, human policy set: $|\Pi|={\pi_{H}}$, agent access budget: $b\sim[0,1]$
  Initialize parameters of $Q_{\theta}$ and $\hat{r}_{\psi}$
  Initialize a dataset of preferences $D_{h}\leftarrow\varnothing$
  // EXPLORATION PHASE
  $D_{\tau},\pi_{\phi}\leftarrow$ EXPLORE() in [15]
  POLICY LEARNING PHASE
  for each iteration do
     // REWARD LEARNING PHASE
     if new episode then
        choose human policy $\pi_{H}$
     end if
     if iteration % $K$ == 0 then
        for m in 1…$M$ do
           $(\sigma^{0},\sigma^{1})\sim$ SAMPLE() in [15]
           Query human partner for $y$
        end for
        for each gradient step do
           Sample mini-batch $\{{(\sigma^{0},\sigma^{1},y)}_{j}\}^{D_{h}}_{j=1}\sim D_{h}$
        end for
        Relabel entire replay buffer $D_{\tau}$ using $\hat{r}_{\psi}$
     end if
     REINFORCEMENT LEARNING PHASE
     for each time-step $t$ do
        if time-step $t\leq$ budget $b$ then
           sample $a_{h}\sim\pi_{H}$
        else
           randomly sample $a_{h}\sim\mathcal{A}^{H}$
        end if
        Collect $s_{t+1}$ by taking $a_{t}\sim\pi_{\phi}(a_{t}|s_{t})$ and $a_{H}$
        Store transitions $D_{\tau}\leftarrow D_{\tau}\{(s_{t},a_{t},{\color[rgb]{1,0,0}a_{h}},s_{t+1},\hat{r}_{\psi}(s_{t}))\}$
     end for
     for each gradient step do
        Sample random mini-batch $\{{\tau_{j}}\}^{D_{\tau}}_{j=1}\sim D_{\tau}$
        Optimize $L^{SAC}_{critic}$ and $L^{SAC}_{actor}$ w.r.t. $\theta$ and $\phi$, respectively [15]
     end for
  end for
A.3 Modified PEBBLE Algorithm for Human Flexibility experiments
Algorithm 3 Adapting Single Agent PbRL for Human-AI Team
  Input: feedback frequency $K$, # queries per feedback session $M$, human policy set: $|\Pi|$, agent access budget: $b={1}$
  Initialize parameters of $Q_{\theta}$ and $\hat{r}_{\psi}$
  Initialize a dataset of preferences $D_{h}\leftarrow\varnothing$
  // EXPLORATION PHASE
  $D_{\tau},\pi_{\phi}\leftarrow$ EXPLORE() in [15]
  POLICY LEARNING PHASE
  for each iteration do
     // REWARD LEARNING PHASE
     if new episode then
        choose human policy $\pi_{H}\in|\Pi|$
     end if
     if iteration % $K$ == 0 then
        for m in 1…$M$ do
           $(\sigma^{0},\sigma^{1})\sim$ SAMPLE() in [15]
           Query human partner for $y$
        end for
        for each gradient step do
           Sample mini-batch $\{{(\sigma^{0},\sigma^{1},y)}_{j}\}^{D_{h}}_{j=1}\sim D_{h}$
        end for
        Relabel entire replay buffer $D_{\tau}$ using $\hat{r}_{\psi}$
     end if
     REINFORCEMENT LEARNING PHASE
     for each time-step $t$ do
        sample $a_{h}\sim\pi_{H}$
        Collect $s_{t+1}$ by taking $a_{t}\sim\pi_{\phi}(a_{t}|s_{t})$ and $a_{H}$
        Store transitions $D_{\tau}\leftarrow D_{\tau}\{(s_{t},a_{t},{\color[rgb]{1,0,0}a_{h}},s_{t+1},\hat{r}_{\psi}(s_{t}))\}$
     end for
     for each gradient step do
        Sample random mini-batch $\{{\tau_{j}}\}^{D_{\tau}}_{j=1}\sim D_{\tau}$
        Optimize $L^{SAC}_{critic}$ and $L^{SAC}_{actor}$ w.r.t. $\theta$ and $\phi$, respectively [15]
     end for
  end for
Table 1: MA MuJoCO - Action Space
Environment Human’s Action Space AI Agent’s Action Space
Walker Right Leg Joints Left Leg Joints
Hopper Thigh Rotor Leg Rotor, Foot Rotor
HalfCheetah Front Rotor Back Rotor
Ant Front Left Leg, Right Back Leg Front Right Leg, Back Leg
Swimmer First Rotor Second Rotor
Appendix B Experiment Details
B.1 State Space and Action Space
We use the available MuJoCo implementation packaged by OpenAI Gym [62] to instantiate our environments. We use the default state space representation given by the package, and the same observation space is shared by both, the human and the AI agents. The details about the action space are given in Table 1.
For MA Highway, we use the Highway environment packages provided by [11] to create our custom multi-agent environments and the associated tasks. We use the continuous action space consisting of (steering, throttle) for both the agent cars.
B.2 Reward Architecture and Embedding Space
Following the implementation of [15, 13, 14] we implement reward model via a neural network and bound the final output using a $\tanh$ activation function : [-1, 1]. For all the experiments, the reward model has three hidden layers with 256 neurons each followed by the prediction layer (with one neuron). We use the ADAM optimizer for training the SAC actor-critic as well as the reward model. The hyperparameters used for PEBBLE, RUNE & SURF are given in Tables 3, 4, and 5.
The input to the reward model is (state, action) tuple. As used previously [14], we stack the state and action vectors and treat them as a single input for the reward model.
B.3 Human Policy & Preference Oracles
We assume an oracle scripted human who uses the underlying reward to correctly provide the binary feedback for their preferences. The feedback is given as follows :
$y(\tau_{0},\tau_{1})=\begin{cases}0,&\text{if}\quad\sum_{i}\tilde{R_{h}}(\tau_{0})>\sum_{i}\tilde{R_{h}}(\tau_{1})\\ 1,&\text{if}\quad\sum_{i}\tilde{R_{h}}(\tau_{0})<\sum_{i}\tilde{R_{h}}(\tau_{1})\\ \end{cases}$ (3)
where, $\tilde{R_{h}}$ is the environment reward being used as the human preference reward function and $\tau_{0},\tau_{1}$ are the queried trajectory pairs. Note that we work under the setting that the preference feedback is binary and therefore if the trajectory returns are equal we uniformly pick a preferred trajectory. This does not pose any problems with our chosen benchmark domains as the underlying reward is dense and shaped [63, 64].
To construct the human policy oracles, we train a SAC agent treating each environment as a single-agent environment, and retrieve the oracle policy for the human’s action space from this SAC agent at the time of training the AI agent.
The hand-crafted rewards for MA Highway - Right Overtake have been shown in Table 2. Similarly, we construct the rewards for MA Highway - Middle & Left Overtake domains.
Table 2: MA Highway - Right Overtake rewards. (x1,y1) is the position of AI agent and (x2,y2) is the position of the Human agent in the environment.
Condition Position Return
Same Lane x1 <x2 -0.5
y1 >= y2 - 2 & y1 <= y2 + 2 x1 >x2 1
Adjacent Lane x1 <= x2 - 10 -0.25
y1 >= 2& y2 <= 6 x1 >= x2 + 10 0.5
else -1
Overtaking Lane x1 <= x2 0.25
y1 >= 6 & y2 <= 10: x1 >= x2 0.25
B.4 Implementation, Code and Compute
We use the publicly available implementation of [15, 13, 14] for the implementation of SAC, PEBBLE, RUNE & SURF algorithms. Our implementation of extending these PbRL algorithms to the multi-agent setting can be found in the code given in the supplementary. All the experiments were run on an Intel(R) Xeon(R) Gold 6258R CPU @ 2.70GHz, with Quadro RTX 8000 GPU.
Appendix C Hyperparameters
Table 3: PEBBLE Hyperparameters
Hyperparameter Value Hyperparameter Value
Initial temperature 0.1 Hidden units per each layer 1024
Segment length 30 (MA Highway) # of layers 2
50 (MA MuJoCo) Optimizer Adam [65]
Learning rate 0.0003 Batch Size 1024(DMControl)
Critic target update freq 2 Critic EMA $\tau$ 0.005
$\beta_{1},\beta_{2}$ (0.9, 0.999) Discount $\gamma$ 0.99
Frequency of feedback 500 (MA Highway) Maximum budget / 800/25 (MA Highway)
10000 (MA MuJoCo) # of queries per session 1400/14 (MA MuJoCo)
# of pre-training steps 1000 (MA Highway)
# of ensemble models $N_{en}$ 3 5000 (MA MuJoCo)
Table 4: SURF Hyperparameters
Hyperparameter Value
Unlabeled batch ratio $\mu$ 4
Threshold $\tau$ 0.95
Segment length 15 (MA Highway)
50 (MA MuJoCo)
Loss weight $\lambda$ 1
Table 5: RUNE Hyperparameters
Hyperparameter Value
Segment length 20 (MA Highway)
50 (MA MuJoCo)